The material to be used for an antireflective coating has index of refreaction of 1.25. How thick should the coating be to give the best result for lambda=550 nm and an angle of incidence of 30 degrees withthe normal
1
Expert's answer
2020-05-18T10:06:29-0400

In our case for destructive interference


"2t\sqrt{n^2-\sin^2i}-\frac{\lambda}{2}=(2k+1)\frac{\lambda}{2} \to"


"2t\sqrt{n^2-\sin^2i}=k\lambda+\frac{\lambda}{2}+\frac{\lambda}{2}=(k+1)\lambda,k=0,1,2,...."


Assume that "k=0"


"2t\sqrt{n^2-\sin^2i}=\lambda\to t=\frac{\lambda}{2\sqrt{n^2-\sin^2i}}="


"=\frac{550\cdot10^{-9}}{2\sqrt{1.25^2-\sin^230°}}=2.4\cdot10^{-7}m"